# Bootstrapping the Abelian Lattice Gauge Theories

### v1

@@ 1. Introduction
The strongly coupled gauge theories are the longstanding problems in theoretical physics. The modern conformal bootstrap program [1, 2] provides a powerful non-perturbative approach to study the conformal phases of these classical theories. The idea is to derive strict constraints on the parameter space of the strongly coupled gauge theories using general consistency conditions combined with the underlying positive structures. The bootstrap implementations are based on the correlation functions of local gauge-invariant operators. Studies along this direction have shown that the bootstrap computations indeed can provide highly non-trivial new results for strongly coupled gauge theories [2], while it also suggests that substantially new ingredients beyond the correlators of local gauge invariant operators are needed to improve the bootstrap results further.
An interesting example for strongly coupled gauge theories is provided by the 3D $U(1)$ gauge theory coupled with $N_{f}$ flavors of two-component Dirac fermions (QED3). QED3 is arguably the simplest theory which provides abundant strongly coupled physics, including the gauge confinement [3, 4, 5], chiral symmetry breaking [6, 7, 8] and conformal phase above the critical flavor number $N_{f}>N_{f}^{*}$ [9, 10]. QED3 has played important roles in the condensed matter physics and has been proposed to have beautiful theoretical properties, such the 3D duality web and symmetry enhancement in the infrared [11]. An important question of QED3 is to determine the critical flavor number $N_{f}^{*}$. The conformal bootstrap results [12, 13, 14], combined with the large $N_{f}$ expansion [15, 16, 17, 18, 19, 20, 21] and the Monte Carlo simulations [22, 23, 24, 25] lead to an estimate for the critical flavor number $N_{f}^{*}\in(2,4)$. Nevertheless, to bootstrap the critical indices of conformal QED3, it needs extra inputs from the large $N_{f}$ perturbative expansions or the lattice simulations, while it is difficult to numerically capture the conformal gauge theories merely from bootstrap constraints. This observation is further supported by the fact that the four-point crossing equations in the bootstrap implementations are endowed with positive algebraic structures, which lead to the bootstrap bound coincidences among different global symmetries [26]. As a consequence, it is hard to obtain bootstrap solutions specific to the gauge theories with various global symmetries and representations.
@@
In summary, the conformal and S-matrix bootstrap studies suggest that the current bootstrap results on strongly coupled gauge theories depend on the extra input data from other approaches, and the bootstrap method is not full-fledged yet to provide solutions for the classical strongly coupled gauge theories. This situation may not be surprising, as both of the bootstrap approaches are based on the gauge invariant local operators, in which the gauge information is not directly specified and details of the gauge dynamics are invisible in the bootstrap setup. Since the nonlocal gauge-invariant operators, e.g., the Wilson loops carry key information on the gauge interactions, it is natural to include the Wilson loops in the bootstrap setup. The long term goal along this direction is to develop a self-contained bootstrap method to study the strongly coupled gauge theories without resorting to additional data. In this work we start with simpler but still physically interesting theories: the Abelian gauge theories with lattice regularization.
@@
In lattice gauge theories, the fundamental physical observables are Wilson loops. Expectation values of the Wilson loops are restricted by the Makeenko-Migdal loop equations [31, 32, 33]. For general gauge theories the loop equations cannot be solved directly. In [34] Anderson and Kruczenski proposed a novel method to bootstrap the loop equations. Their key observation is that the matrices made from the inner products of the Wilson lines are positive definite. Therefore the Wilson loop averages are restricted by both the loop equations and the positive definiteness of these matrices. It turns out that the two conditions together can lead to strong constraints on the Wilson loop averages. This approach has been further improved in a recent work by Kazakov and Zheng [35], in which the reflection positivity [36] in the lattice gauge theories has been employed for bootstrap computations. The authors also introduced the relaxation procedure to deal with the nonlinear parts in the loop equations. The lattice gauge theory bootstrap is closely related to a series of bootstrap works on matrix models [37, 38, 39, 40, 41], quantum mechanics [42, 43, 44, 45, 46, 47, 48], many body systems and the Ising model [49, 50], etc. [^1] See also the bootstrap studies of classical system [51] and null state bootstrap without imposing positivity [52, 53]
@@
We study the $\mathbb{Z}_{2}$ and $U(1)$ Abelian lattice gauge theories. The two theories can be solved analytically in 2D. The 3D $\mathbb{Z}_{2}$ lattice gauge theory is dual to the 3D Ising model, and provides a 3D generalization of the Kramers-Wannier duality [54]. It is the first known example for phase transitions without local order parameter and provides a useful model for the quantum computing [55] and generalized global symmetry [56]. The 3D $U(1)$ lattice gauge theory is one of few examples for which the confinement mechanism can be understood analytically, and it plays an instructive role in the studies of gauge confinement. Semi-classical analysis suggests that the 3D $U(1)$ lattice gauge theory is confined in the long-distance limit for any gauge coupling [3, 4, 5], and the confinement is triggered by condensation of the $U(1)$ monopoles. This theory has been extensively studied since 1970s by different approaches in order to gain detailed understanding of the confinement; see recent Monte Carlo simulations to extract the string tension and glueball spectrum [57, 58, 59]. From our point view, the Abelian lattice gauge theories provide natural candidates to explore the bootstrap constraints from the non-local gauge invariant operators.
@@
In this work, the first question we want to address is whether the positivity conditions and the loop equations can have sufficiently strong constraints to numerically solve the lattice gauge theories, in the sense of computing the parameters of the theories to high numerical precision. The previous two-sided bootstrap bounds are rather strong but not enough to extract precise data on the confinement, like the string tension and glueball spectrum. A technical obstacle is that to bootstrap the loop equations, it needs to truncate the length of the Wilson loops with a cutoff $L_{\textrm{max}}$, and the numerical results show strong dependence on the cutoff $L_{\textrm{max}}$ up to $L_{\textrm{max}}=16$ [34, 35]. [^2] Such dependence of the bootstrap bounds on the size of operators has also been shown in the bootstrap results of the Ising model on the lattice [49]. One may expect the bootstrap results with much higher cutoff $L_{\textrm{max}}$ could improve the results significantly. Nevertheless, the computational resource needed for the bootstrap studies increases drastically and quickly becomes impractical with higher $L_{\textrm{max}}$. [^3] See [60] for new bootstrap results with $L_{\textrm{max}}=20,24$. We find that this problem can be nicely resolved in bootstrapping 2D $U(1)$ lattice gauge theory by fixing the gauge redundancy. The two-sided bootstrap bounds can give numerical estimates of the Wilson loop averages to numerical precision at $10^{-8}$ or higher with $L_{\textrm{max}}=60$, thus providing strong evidence that the loop equations and the positivity conditions contain sufficient strong constraints to numerically solve this lattice gauge theory.
@@
In 3D the gauge fixing does not simplify the loop equations, and we apply the bootstrap method developed in [34, 35] for the $\mathbb{Z}_{2}$ and $U(1)$ lattice gauge theories. In [34, 35] the authors focused on the $SU(N)$ lattice gauge theories in the large $N$ limit and there are two notable differences for the Abelian lattice gauge theories studied in this work. Firstly, different from the Yang-Mills lattice gauge theories studied in [34, 35], the loop equations for the Abelian lattice gauge theories are linear so no need to introduce the relaxation procedure. Besides, the real sectors of the Abelian lattice gauge theories are known to satisfy Griffiths’ inequalities [61, 62, 63]. The bootstrap study of the Griffiths’ inequalities was firstly introduced in [49]. The Griffiths’ inequalities can be used to prove the convergence of the lattice system in the infinite volume limit. They have been employed to derive various interesting bounds on the lattice theories nearly half a century before the modern bootstrap endeavors. We will show that they have interesting connections with our bootstrap bounds.
@@ 3.2.  Griffiths’ inequalities in Abelian lattice gauge theories
Besides the Ising model, the Griffiths’ inequalities have been known to hold in many lattice models [67, 62, 63]. Especially in [63], Ginibre has proved the Griffiths’ inequalities in sufficient generality. Here we explain the Ginibre’s proof related to the Abelian lattice gauge theories. We clarify a subtlety due to which the inequalities are not always satisfied, and for some cases, even the direction of the inequality can be switched.
@@
$\cos{n_{\ell}\theta_{\ell}}\pm\cos{n_{\ell}\theta^{\prime}_{\ell}}$ $=2\cos{n_{\ell}\frac{\theta_{\ell}\mp\theta^{\prime}_{\ell}}{2}}\cos{n_{\ell}\frac{\theta_{\ell}\pm\theta^{\prime}_{\ell}}{2}},$ (3.14)
@@
3.3.  A universal lower bound from Griffiths’ inequality
@@ 4.2.  Bootstrap bounds on the 2D $U(1)$ lattice gauge theory
We have computed the bootstrap bounds up to $L_{\textrm{max}}=60$ and studied the precision of bootstrap bounds with increasing $L_{\textrm{max}}$. The results are shown in Fig. 2. In the left panel, we show the precision of the two-sided bounds for different $\lambda$. The two-sided bounds converge much better in the strong and weakly coupled regions. The precision becomes weaker for the intermediate $\lambda$. In the right panel of Fig. 2, we show the maximum spread, or the precision of the two-sided bounds for fixed $L_{\textrm{max}}$. The precision $\delta w_{1}$ improves linearly with increasing $L_{\textrm{max}}$. At $L_{\textrm{max}}=60$ it reaches a remarkably high precision $10^{-8}$!  [^9] The bounds in Figs. 1 and 2 are computed using the “SemidefiniteOptimization” program in Mathematica, with “Tolerance” $10^{-8}$. The precision at $L_{\textrm{max}}=60$ is close to the “Tolerance” value so the results are partially affected by the numerical setup. It clearly shows that the precision of the two-sided bootstrap bounds is proportional to the maximum length of the Wilson loops $L_{\textrm{max}}$: $-\ln(\delta w_{1})\propto L_{\textrm{max}}$.
@@ 5.2.  Bootstrapping the 3D $\mathbb{Z}_{2}$ lattice gauge theories
which is represented by the red dot-dashed line in the figure. In the strong coupling limit (small $J$), the two-sided bootstrap bounds indeed coincide with the perturbative result with high precision. Another interesting property is that for $J<0.6$, the lower bound is close to the exact 2D solution of the $\mathbb{Z}_{2}$ lattice gauge theory: $u_{P}=\tanh{J}$, given by the green dot-dashed line in the figure. In fact, with $L_{\textrm{max}}=12$, the lower bound on $u_{P}$ almost coincides with the exact 2D solution for $J<0.62$. This confirms again the robustness of the lower bound on $u_{P}$ from the Griffiths’ inequalities (3.27), and the modern bootstrap algorithm can move much further.
Figure 5: Bootstrap bounds on the plaquette average $u_{P}$ in the 3D $\mathbb{Z}_{2}$ lattice gauge theory. The light (dark) shadowed region denotes the bootstrap allowed region for $u_{P}$ with maximum Wilson loop length $L_{\textrm{max}}=12$ ($L_{\textrm{max}}=16$). The red dot-dashed line represents the strong coupling expansion of $u_{P}$ to the order $J^{15}$. Solution of the 2D $\mathbb{Z}_{2}$ lattice gauge theory $u_{P}=\tanh{J}$ is denoted by the green dot-dashed line. The purple vertical dashed line corresponds to $J^{*}\approx 0.761413$, where the confinement-deconfinement phase transition happens.
Different from the 3D $U(1)$ lattice gauge theory, the 3D $\mathbb{Z}_{2}$ lattice gauge theory admits a second order phase transition corresponding to the deconfinement process, which is dual to the phase transition in the 3D Ising model. The critical coupling coefficient $J^{\prime}$ for the 3D Ising model has been estimated to extremely high precision using large-scale Monte Carlo simulation [77]:
$J^{\prime}=0.221654626(5).$ (5.5)

### v2

@@ 1. Introduction
The strongly coupled gauge theories are the longstanding problems in theoretical physics. The modern conformal bootstrap program [1, 2] provides a powerful non-perturbative approach to study the conformal phases of these classical theories. The idea is to derive strict constraints on the parameter space of the strongly coupled gauge theories using general consistency conditions combined with the underlying positive structures. The bootstrap implementations are based on the correlation functions of local gauge-invariant operators. Studies along this direction have shown that the bootstrap computations indeed can provide highly non-trivial new results for strongly coupled gauge theories [2], while it is also suggested that substantially new ingredients beyond the correlators of local gauge invariant operators are needed to improve the bootstrap results further.
An interesting example for strongly coupled gauge theories is provided by the 3D $U(1)$ gauge theory coupled with $N_{f}$ flavors of two-component Dirac fermions (QED3). QED3 is arguably the simplest theory which provides abundant strongly coupled physics, including the gauge confinement [3, 4, 5], chiral symmetry breaking [6, 7, 8] and conformal phase above the critical flavor number $N_{f}>N_{f}^{*}$ [9, 10]. QED3 has played important roles in the condensed matter physics and has been proposed to have beautiful theoretical properties, such as the 3D duality web and symmetry enhancement in the infrared [11]. An important question of QED3 is to determine the critical flavor number $N_{f}^{*}$. The conformal bootstrap results [12, 13, 14], combined with the large $N_{f}$ expansion [15, 16, 17, 18, 19, 20, 21] and the Monte Carlo simulations [22, 23, 24, 25] lead to an estimate for the critical flavor number $N_{f}^{*}\in(2,4)$. Nevertheless, to bootstrap the critical indices of conformal QED3, it needs extra inputs from the large $N_{f}$ perturbative expansions or the lattice simulations, while it is difficult to numerically capture the conformal gauge theories merely from bootstrap constraints. This observation is further supported by the fact that the four-point crossing equations in the bootstrap implementations are endowed with positive algebraic structures [26, 27], which lead to the bootstrap bound coincidences among different global symmetries [28]. As a consequence, it is hard to obtain bootstrap solutions specific to the gauge theories with various global symmetries and representations.
@@
In summary, the conformal and S-matrix bootstrap studies suggest that the current bootstrap results on strongly coupled gauge theories depend on the extra input data from other approaches, and the bootstrap method is not full-fledged yet to provide solutions for the classical strongly coupled gauge theories. This situation may not be surprising, as both of the bootstrap approaches are based on the gauge invariant local operators, in which the gauge information is not directly specified and details of the gauge interactions are invisible in the bootstrap setup. Since the nonlocal gauge-invariant operators, e.g., the Wilson loops carry key information on the gauge interactions, it is natural to include the Wilson loops in the bootstrap setup. The long term goal along this direction is to develop a self-contained bootstrap method to study the strongly coupled gauge theories without resorting to additional data. In this work we start with simpler but still physically interesting theories: the Abelian gauge theories with lattice regularization.
@@
In lattice gauge theories, the fundamental physical observables are Wilson loops. Expectation values of the Wilson loops are restricted by the Makeenko-Migdal loop equations [33, 34, 35]. For general gauge theories the loop equations cannot be solved directly. In [36] Anderson and Kruczenski proposed a novel method to bootstrap the loop equations. Their key observation is that the matrices made from the inner products of the Wilson lines are positive definite. Therefore the Wilson loop averages are restricted by both the loop equations and the positive definiteness of these matrices. It turns out that the two conditions together can lead to strong constraints on the Wilson loop averages. This approach has been further improved in a recent work by Kazakov and Zheng [37], in which the reflection positivity [38] in the lattice gauge theories has been employed for bootstrap computations. The authors also introduced the relaxation procedure to deal with the nonlinear parts in the loop equations. The lattice gauge theory bootstrap is closely related to a series of bootstrap works on matrix models [39, 40, 41, 42, 43], quantum mechanics [44, 45, 46, 47, 48, 49, 50, 51], many body systems [52, 53] and the Ising model [54], etc. [^1] See also the bootstrap studies of classical system [55] and null state bootstrap without imposing positivity [56, 57]
@@
We study the $\mathbb{Z}_{2}$ and $U(1)$ Abelian lattice gauge theories. The two theories can be solved analytically in 2D. The 3D $\mathbb{Z}_{2}$ lattice gauge theory is dual to the 3D Ising model, and provides a 3D generalization of the Kramers-Wannier duality [58]. It is the first known example for phase transitions without local order parameter and provides a useful model for the quantum computing [59] and generalized global symmetry [60]. The 3D $U(1)$ lattice gauge theory is one of few examples for which the confinement mechanism can be understood analytically, and it plays an instructive role in the studies of gauge confinement. Semi-classical analysis suggests that the 3D $U(1)$ lattice gauge theory is confined in the long-distance limit for any gauge coupling [3, 4, 5], and the confinement is triggered by condensation of the $U(1)$ monopoles. This theory has been extensively studied since 1970s by different approaches in order to gain detailed understanding of the confinement, see recent Monte Carlo simulations to extract the string tension and glueball spectrum [61, 62, 63]. From our point of view, the Abelian lattice gauge theories provide natural candidates to explore the bootstrap constraints from the non-local gauge invariant operators.
@@
In this work, the first question we want to address is whether the positivity conditions and the loop equations have sufficiently strong constraints to numerically solve the lattice gauge theories, in the sense of computing the parameters of the theories to high numerical precision. The previous two-sided bootstrap bounds are rather impressive but not enough to extract precise data on the confinement, like the string tension and glueball spectrum. A technical obstacle is that to bootstrap the loop equations, it needs to truncate the length of the Wilson loops with a cutoff $L_{\textrm{max}}$, and the numerical results show strong dependence on the cutoff $L_{\textrm{max}}$ up to $L_{\textrm{max}}=16$ [36, 37]. [^2] Such dependence of the bootstrap bounds on the truncation of operators has also been shown in the bootstrap results of the Ising model on the lattice [54]. One may expect the bootstrap results with much higher cutoff $L_{\textrm{max}}$ could improve the results significantly. Nevertheless, the computational resource needed for the bootstrap studies increases drastically and quickly becomes impractical with higher $L_{\textrm{max}}$. [^3] See [64] for new bootstrap results with $L_{\textrm{max}}=20,24$. We find that this problem can be nicely resolved in bootstrapping 2D $U(1)$ lattice gauge theory by fixing the gauge redundancy. The two-sided bootstrap bounds can give numerical estimates of the Wilson loop averages to numerical precision near $10^{-8}$ or even higher with $L_{\textrm{max}}=60$, thus providing strong evidence that the loop equations and the positivity conditions contain sufficiently strong constraints to numerically solve this lattice gauge theory.
@@
In 3D the gauge fixing does not simplify the loop equations, and we apply the bootstrap method developed in [36, 37] for the $\mathbb{Z}_{2}$ and $U(1)$ lattice gauge theories. In [36, 37] the authors focused on the $SU(N)$ lattice gauge theories in the large $N$ limit and there are two notable differences for the Abelian lattice gauge theories studied in this work. Firstly, different from the Yang-Mills lattice gauge theories studied in [36, 37], the loop equations for the Abelian lattice gauge theories are linear so no need to introduce the relaxation procedure. Besides, the real sectors of the Abelian lattice gauge theories satisfy Griffiths’ inequalities [65, 66, 67]. The bootstrap study of the Griffiths’ inequalities was firstly introduced in [54]. The Griffiths’ inequalities can be used to prove the convergence of the lattice system in the infinite volume limit. They have been employed to derive various interesting bounds on the lattice theories nearly half a century before the modern bootstrap endeavors. We will show that they have interesting connections with our bootstrap bounds.
@@ 3.2.  Griffiths’ inequalities in Abelian lattice gauge theories
Besides the Ising model, the Griffiths’ inequalities have been known to hold in many lattice models [71, 66, 67]. Especially in [67], Ginibre has proved the Griffiths’ inequalities in sufficient generality. Here we explain Ginibre’s proof related to the Abelian lattice gauge theories. We clarify a subtlety due to which the inequalities are not always satisfied, and for some cases, even the direction of the inequality can be switched.
@@
$\begin{split}\cos{n_{\ell}\theta_{\ell}}+\cos{n_{\ell}\theta^{\prime}_{\ell}}&=2\cos{n_{\ell}\frac{\theta_{\ell}-\theta^{\prime}_{\ell}}{2}}\cos{n_{\ell}\frac{\theta_{\ell}+\theta^{\prime}_{\ell}}{2}},\\ \cos{n_{\ell}\theta_{\ell}}-\cos{n_{\ell}\theta^{\prime}_{\ell}}&=2\sin{n_{\ell}\frac{\theta_{\ell}-\theta^{\prime}_{\ell}}{2}}\sin{n_{\ell}\frac{\theta_{\ell}+\theta^{\prime}_{\ell}}{2}},\end{split}$ (3.14)
@@
3.3.  A universal lower bound from Griffiths’ inequalities
@@ 4.2.  Bootstrap bounds on the 2D $U(1)$ lattice gauge theory
Table 1: The first 10 digits of the bootstrap bounds on the averaged plaquette value $w_{1}$ in 2D $U(1)$ lattice gauge theory with $L_{\textrm{max}}=60$.
   $\lambda$      Lower bound      Upper bound      Exact value
$0.01$ $0.99749685{\color[rgb]{1,0,0}14}$ $0.997496859{\color[rgb]{1,0,0}5}$ $0.9974968592$
$0.1$ $0.97467050{\color[rgb]{1,0,0}40}$ $0.974670517{\color[rgb]{1,0,0}2}$ $0.9746705078$
$0.2$ $0.9485998{\color[rgb]{1,0,0}188}$ $0.9485998{\color[rgb]{1,0,0}342}$ $0.9485998259$
$0.4$ $0.89338313{\color[rgb]{1,0,0}27}$ $0.8933831{\color[rgb]{1,0,0}412}$ $0.8933831370$
$0.6$ $0.83190007{\color[rgb]{1,0,0}05}$ $0.831900071{\color[rgb]{1,0,0}7}$ $0.8319000711$
$0.8$ $0.764996747{\color[rgb]{1,0,0}0}$ $0.76499674{\color[rgb]{1,0,0}81}$ $0.7649967475$
$1.0$ $0.697774657{\color[rgb]{1,0,0}5}$ $0.69777465{\color[rgb]{1,0,0}83}$ $0.6977746579$
$2.0$ $0.446389965{\color[rgb]{1,0,0}6}$ $0.44638996{\color[rgb]{1,0,0}61}$ $0.4463899659$
$4.0$ $0.2424996125$ $0.242499612{\color[rgb]{1,0,0}6}$ $0.2424996125$
$6.0$ $0.1643939155$ $0.1643939155$ $0.1643939155$
Refer to captionRefer to caption
@@
We have computed the bootstrap bounds up to $L_{\textrm{max}}=60$ and studied the precision of bootstrap bounds with increasing $L_{\textrm{max}}$. The first ten digits of the upper and lower bounds with $L_{\textrm{max}}=60$ are shown in Table. 1, in which the numerical precision is of order $10^{-8}$ for $\lambda\sim 0.2$ and becomes even better for other $\lambda$s. The numerical precision for different $L_{\textrm{max}}$ is shown in Fig. 2. In the left panel, we show the precision of the two-sided bounds for different $\lambda$. The two-sided bounds converge much better in the strong and weakly coupled regions. The precision becomes weaker for the intermediate $\lambda$. In the right panel of Fig. 2, we show the maximum spread, or the precision of the two-sided bounds for fixed $L_{\textrm{max}}$. The precision $\delta w_{1}$ improves linearly with increasing $L_{\textrm{max}}$. At $L_{\textrm{max}}=60$ it reaches a remarkably high precision $10^{-8}$!  [^9] The bounds in Figs. 1 and 2 are computed using the “SemidefiniteOptimization” program in Mathematica, with “Tolerance” $10^{-8}$. The precision at $L_{\textrm{max}}=60$ is close to the “Tolerance” value so the results are partially affected by the numerical setup. It clearly shows that the precision of the two-sided bootstrap bounds is proportional to the maximum length of the Wilson loops $L_{\textrm{max}}$: $-\ln(\delta w_{1})\propto L_{\textrm{max}}$.
@@ 5.2.  Bootstrapping the 3D $\mathbb{Z}_{2}$ lattice gauge theories
which is represented by the red dashed line in the figure. In the strong coupling limit (small $J$), the two-sided bootstrap bounds indeed coincide with the perturbative result with high precision. Another interesting property is that for $J<0.6$, the lower bound is close to the exact 2D solution of the $\mathbb{Z}_{2}$ lattice gauge theory: $u_{P}=\tanh{J}$, given by the black dashed line in the figure. In fact, with $L_{\textrm{max}}=12$, the lower bound on $u_{P}$ almost coincides with the exact 2D solution for $J<0.62$. This confirms again the robustness of the lower bound on $u_{P}$ from the Griffiths’ inequalities (3.27), and the modern bootstrap algorithm can move much further.
Figure 5: Bootstrap bounds on the plaquette average $u_{P}$ in the 3D $\mathbb{Z}_{2}$ lattice gauge theory. The light (dark) blue shadowed region denotes the bootstrap allowed region for $u_{P}$ with maximum Wilson loop length $L_{\textrm{max}}=12$ ($L_{\textrm{max}}=16$). The green dots are from Mento Carlo data for the 3D Ising model [54], which has been transferred to the 3D $\mathbb{Z}_{2}$ lattice gauge theory through the duality relation (5.5). The red dashed line represents the strong coupling expansion of $u_{P}$ to the order $J^{15}$. Solution of the 2D $\mathbb{Z}_{2}$ lattice gauge theory $u_{P}=\tanh{J}$ is denoted by the black dashed line. The purple vertical dashed line corresponds to $J=J^{*}\approx 0.761413$, where the confinement-deconfinement phase transition happens.
Different from the 3D $U(1)$ lattice gauge theory, the 3D $\mathbb{Z}_{2}$ lattice gauge theory admits a second order phase transition corresponding to the deconfinement process. It is well known that the 3D $\mathbb{Z}_{2}$ lattice gauge theory is dual to the 3D Ising model. Denote the coupling coefficient of the dual 3D Ising model by $J^{\prime}$, the free energies of the two theories satisfy
$F_{\textrm{Ising}}(J^{\prime})=F_{\mathbb{Z}_{2}\textrm{ gauge}}(J)-\frac{3}{2}\ln\sinh(2J)+\frac{1}{2}\ln 2.$ (5.5)
By taking derivatives with respect to $J/J^{\prime}$, above duality relation leads to an equation between the plaquette average $u_{P}$ and the nearest-spin correlator $\langle s_{i}s_{i+\mu}\rangle$ in the 3D Ising model. We use the Monte Carlo data on the nearest-spin correlator $\langle s_{i}s_{i+\mu}\rangle$ presented in [54] to compute the plaquette average $u_{P}$. The results are given by the green dots in Fig. 5, which show excellent agreements with the bootstrap bounds with $L_{\textrm{max}}=16$. The critical coupling coefficient ${J^{\prime}}^{*}$ for the 3D Ising model has been estimated to extremely high precision using large-scale Monte Carlo simulation [81]:
${J^{\prime}}^{*}=0.221654626(5).$ (5.6)